if six numbers are picked from 1 to 45 and if a number is not repeated.How many sets of six numbers could be picked
Answer by jim_thompson5910(35256)

I'm assuming order does not matter.

C(n, r) = (n!)/(r!(n-r)!)

C(45, 6) = (45!)/(6!*(45-6)!)

C(45, 6) = (45!)/(6!*39!)

C(45, 6) = (45*44*43*42*41*40*39!)/(6!*39!)

C(45, 6) = (45*44*43*42*41*40)/(6!)

C(45, 6) = (45*44*43*42*41*40)/(6*5*4*3*2*1)

C(45, 6) = (5864443200)/(720)

C(45, 6) = 8145060

So there are 8145060 different ways.